# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@ 5 Simulation
where $u_{t}=(u_{t,1},\cdots,u_{t,r_{0}})^{{}^{\prime}}$ is i.i.d. $N(0,I_{r_{0}})$ for $t=2,\cdots,T$ and $f_{1}=(f_{1,1},\cdots,f_{1,r_{0}})^{{}^{\prime}}$ is i.i.d. $N(0,\frac{1}{1-\rho^{2}}I_{r_{0}})$. The scalar $\rho$ captures the serial correlation of factors, and the idiosyncratic errors are generated by
@@
where $v_{t}=(v_{1,t},\cdots,v_{N,t})^{{}^{\prime}}$ is i.i.d. $N(0,\Omega)$ for $t=2,\cdots,T$ and $e_{1}=(e_{1,1},\cdots,e_{N,1})^{{}^{\prime}}$ is $N(0,\frac{1}{(1-\alpha^{2})}\Omega)$. The scalar $\alpha$ captures the serial correlation of the idiosyncratic errors, and $\Omega$ is generated as $\Omega_{ij}=\beta^{|i-j|}$ so that $\beta$ captures the degree of cross-sectional dependence of the idiosyncratic errors. In addition, $u_{t}$ and $v_{t}$ are mutually independent for all values of $t$. We consider the following DGPs for factor loadings and investigate the performance of the QML estimator for different types of breaks discussed in Section 3.
@@
DGP 1.A In this setup, we adopt the DGP from MS. Each of the three regimes has $r_{0}$ effective factors, i.e., $r_{1}=r_{2}=r_{3}=3$. The factor loading matrices on the effective factors are generated as:
@@
The $\Lambda$ matrix in (17) can be written as
$\Lambda\;=\;\bigl{[}\breve{\Lambda}_{1},\breve{\Lambda}_{2},\breve{\Lambda}_{3}\bigr{]}_{N\times 3r_{0}}.$
@@
Define the three $3r_{0}\times 3r_{0}$ block-selector matrices
$B_{1}=\begin{pmatrix}I_{r_{0}}&0&0\\ 0&0_{r_{0}\times r_{0}}&0\\ 0&0&0_{r_{0}\times r_{0}}\end{pmatrix},\quad B_{2}=\begin{pmatrix}0_{r_{0}\times r_{0}}&0&0\\ 0&I_{r_{0}}&0\\ 0&0&0_{r_{0}\times r_{0}}\end{pmatrix},\quad B_{3}=\begin{pmatrix}0_{r_{0}\times r_{0}}&0&0\\ 0&0_{r_{0}\times r_{0}}&0\\ 0&0&I_{r_{0}}\end{pmatrix}$,
so that in regime $j$ we set
$\Lambda_{j}=\Lambda\,B_{j},$
@@
which yields three effective factors per regime and a total of $r=3r_{0}=9$ pseudo-factors over the full sample. This corresponds to a Type 1 breakpoint scenario with $r_{j,j+1}=r_{j}+r_{j+1}$, where loadings in adjacent regimes are independent. By varying $b$ we can control the magnitude of the break.
Since the computation of $\hat{k}_{MS}$ requires the number of original factors ($r_{0}$), we use $\hat{r}=r_{0}$ for it. Moreover, as $\hat{k}_{QML}$ and $\hat{k}_{BKW}$ require the number of pseudo-factors ($r$), we set $\hat{r}=r$ for them. This allows us to focus on the estimation errors of the breakpoint estimators themselves, rather than errors resulting from misspecifying the number of factors.
Table 1 report the RMSEs and MAEs of the breakpoint estimators $\hat{k}_{\mathrm{BKW}}$, $\hat{k}_{\mathrm{MS}}$, and $\hat{k}_{\mathrm{QML}}$ for $(\rho,\alpha,\beta)$ under $b=1$ and $b=0$. We see that even for small samples, the QML estimator’s RMSE and MAE are nearly zero, and they converge to zero as $(N,T)$ grow, confirming the consistency result in Theorem 2(i). The BKW and MS estimators also perform well but remain less accurate than QML.
Table 1: Simulated root mean squared errors (RMSEs) and mean absolute errors (MAEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, under DGP 1.A with $b=1$ (left) and $b=0$ (right).
$b=1$ $b=0$
QML BKW MS QML BKW MS
$N,T$ RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
$\rho=0,\alpha=0,\beta=0$
100,100 (0.148,0.134) (0.022,0.018) (1.166,1.261) (0.356,0.335) (1.396,6.403) (0.550,1.500) (0.182,0.195) (0.033,0.036) (1.017,0.903) (0.285,0.299) (1.265,0.894) (0.800,0.400)
100,300 (0.110,0.148) (0.012,0.018) (0.586,0.688) (0.205,0.233) (0.427,0.334) (0.097,0.063) (0.173,0.158) (0.22,0.023) (0.706,0.621) (0.245,0.225) (2.652,2.956) (0.417,0.411)
300,300 (0.055,0.055) (0.003,0.003) (0.696,0.638) (0.243,0.215) (0.554,0.502) (0.156,0.126) (0.055,0.078) (0.003,0.006) (0.691,0.575) (0.228,0.196) (0.951,0.787) (0.453,0.312)
300,600 (0.055,0.045) (0.003,0.002) (0.625,0.612) (0.215,0.208) (0.685,0.653) (0.236,0.213) (0.063,0.045) (0.004,0.002) (0.519,0.459) (0.175,0.153) (1.177,0.923) (0.693,0.425)
600,600 (0.000,0.032) (0.000,0.001) (0.611,0.538) (0.213,0.185) (0.825,0.687) (0.340,0.236) (0.045,0.045) (0.002,0.002) (0.533,0.544) (0.178,0.170) (1.270,0.939) (0.806,0.441)
$\rho=0.7,\alpha=0,\beta=0$
100,100 (0.078,0.078) (0.006,0.006) (2.675,3.631) (0.702,0.926) (2.894,6.358) (0.837,1.767) (0.089,0.114) (0.008,0.011) (2.963,2.075) (0.765,0.492) (3.691,1.581) (2.375,1.250)
100,300 (0.063,0.071) (0.004,0.005) (0.987,1.883) (0.347,0.455) (0.732,0.590) (0.270,0.176) (0.095,0.055) (0.009,0.003) (1.169,1.752) (0.338,0.397) (0.931,0.805) (0.432,0.326)
300,300 (0.000,0.032) (0.000,0.001) (1.569,1.072) (0.329,0.341) (0.902,0.660) (0.407,0.218) (0.032,0.000) (0.001,0.000) (1.238,1.157) (0.364,0.312) (1.060,0.922) (0.561,0.425)
300,600 (0.045,0.045) (0.002,0.002) (0.808,1.002) (0.258,0.303) (0.967,0.769) (0.466,0.297) (0.032,0.045) (0.001,0.002) (0.865,0.909) (0.265,0.260) (1.350,0.934) (0.912,0.438)
600,600 (0.000,0.032) (0.000,0.001) (1.084,1.079) (0.368,0.337) (1.061,0.867) (0.563,0.377) (0.000,0.045) (0.000,0.002) (0.932,0.922) (0.275,0.294) (1.257,0.971) (0.790,0.471)
$\rho=0,\alpha=0.3,\beta=0$
100,100 (0.170,0.164) (0.025,0.025) (0.921,1.165) (0.300,0.321) (0.697,0.775) (0.257,0.314) (0.251,0.205) (0.057,0.040) (1.041,1.194) (0.303,0.322) (4.000,5.344) (1.333,3.000)
100,300 (0.138,0.114) (0.019,0.13) (0.636,0.748) (0.231,0.257) (0.431,0.324) (0.102,0.059) (0.148,0.155) (0.022,0.022) (0.676,0.659) (0.241,0.226) (3.437,1.814) (0.482,0.389)
300,300 (0.055,0.000) (0.003,0.000) (0.665,0.683) (0.228,0.248) (0.600,0.516) (0.182,0.134) (0.063,0.071) (0.004,0.005) (0.562,0.580) (0.190,0.206) (1.082,0.781) (0.586,0.302)
300,600 (0.032,0.045) (0.001,0.002) (0.636,0.528) (0.217,0.177) (0.687,0.661) (0.237,0.217) (0.055,0.063) (0.003,0.004) (0.509,0.465) (0.169,0.146) (1.196,0.958) (0.715,0.463)
600,600 (0.032,0.045) (0.001,0.002) (0.528,0.595) (0.187,0.188) (0.863,0.742) (0.373,0.275) (0.055,0.045) (0.003,0.002) (0.501,0.506) (0.175,0.150) (1.294,0.909) (0.838,0.408)
$\rho=0,\alpha=0,\beta=0.3$
100,100 (0.145,0.148) (0.019,0.022) (1.318,1.157) (0.338,0.339) (3.919,6.342) (1.167,1.262) (0.179,0.192) (0.032,0.033) (1.102,1.111) (0.287,0.286) (1.069,0.756) (0.571,0.286)
100,300 (0.100,0.118) (0.010,0.014) (0.660,0.842) (0.212,0.286) (0.381,0.404) (0.077,0.088) (0.145,0.176) (0.021,0.031) (0.622,0.656) (0.197,0.218) (5.322,3.584) (0.561,0.521)
300,300 (0.078,0.055) (0.006,0.003) (0.593,0.642) (0.211,0.210) (0.547,0.522) (0.152,0.137) (0.055,0.045) (0.003,0.002) (0.546,0.610) (0.182,0.194) (1.024,0.836) (0.526,0.352)
300,600 (0.105,0.032) (0.011,0.001) (0.623,0.573) (0.202,0.196) (0.772,0.648) (0.302,0.210) (0.055,0.045) (0.003,0.002) (0.496,0.431) (0.162,0.134) (1.203,1.023) (0.724,0.525)
600,600 (0.032,0.000) (0.001,0.000) (0.569,0.536) (0.198,0.191) (0.818,0.678) (0.334,0.230) (0.055,0.045) (0.003,0.002) (0.507,0.503) (0.175,0.149) (1.255,0.893) (0.788,0.401)
$\rho=0.7,\alpha=0.3,\beta=0.3$
100,100 (0.118,0.095) (0.012,0.009) (2.471,3.445) (0.650,0.873) (2.121,5.266) (0.750,1.075) (0.100,0.114) (0.008,0.011) (3.115,2.989) (0.788,0.718) (4.336,1.265) (2.800,0.800)
100,300 (0.055,0.055) (0.003,0.003) (2.317,1.338) (0.418,0.409) (0.744,0.532) (0.278,0.143) (0.089,0.141) (0.008,0.012) (1.159,0.967) (0.342,0.298) (0.933,0.880) (0.436,0.389)
300,300 (0.000,0.055) (0.000,0.003) (1.167,1.200) (0.364,0.372) (0.969,0.693) (0.469,0.241) (0.000,0.055) (0.000,0.003) (1.099,2.137) (0.366,0.328) (1.153,0.920) (0.665,0.424)
300,600 (0.045,0.032) (0.002,0.001) (1.046,1.051) (0.358,0.335) (0.985,0.823) (0.486,0.339) (0.032,0.032) (0.001,0.001) (0.921,1.050) (0.269,0.304) (1.298,0.943) (0.843,0.444)
600,600 (0.032,0.032) (0.001,0.001) (0.818,1.030) (0.271,0.319) (1.044,0.870) (0.543,0.379) (0.032,0.000) (0.001,0.000) (0.788,0.980) (0.237,0.255) (1.429,0.916) (1.021,0.420)
@@
DGP 1.B We set $\lambda_{i}$ to be i.i.d. $N(0,\frac{1}{r}I_{r})$ across $i$, and define $\Lambda=(\lambda_{1},\ldots,\lambda_{N})^{\prime}$. In the first regime, $\Lambda_{1}=(\lambda_{1,1},\cdots,\lambda_{N,1})^{{}^{\prime}}=\Lambda B_{1}$ with $B_{1}=[1,0,0;0,1,0;0,0,0]$. In the second regime, $\Lambda_{2}=(\lambda_{1,2},\cdots,\lambda_{N,2})^{{}^{\prime}}=\Lambda B_{2}$ with $B_{2}=[1,0,0;0,0,0;0,0,1]$. In the third regime, $\Lambda_{3}=(\lambda_{1,3},\cdots,\lambda_{N,3})^{{}^{\prime}}=\Lambda B_{3}$ with $B_{3}=[0,0,0;0,1,0;0,0,1]$. For the full sample, the number of pseudo-factors is $r=3$. In each of the three regimes, the number of efffective factors is 2, as $\text{rank}(B_{1})=\text{rank}(B_{2})=\text{rank}(B_{3})=2$. This setting corresponds to a Type 1 break in Section 3 with $r_{j,j+1}>\max\{r_{j},r_{j+1}\}$. In this DGP, each regime has an inactive factor. Table 2 lists the RMSEs and MAEs of these breakpoint estimators $\hat{k}_{BKW}$, $\hat{k}_{MS}$, and $\hat{k}_{QML}$ with different values of $(\rho,\alpha,\beta)$.
In all cases, $\hat{k}_{QML}$ has much smaller RMSEs and MAEs than $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, and the RMSEs and MAEs of $\hat{k}_{QML}$ tend to decrease as $N$ and $T$ increase. This confirms the consistency of $\hat{k}_{QML}$ established in Theorem 2 (i). In addition, the RMSEs and MAEs of $\hat{k}_{BKW}$ do not converge to zero as $N$ and $T$ increase, indicating that the estimation error of $\hat{k}_{BKW}$ remains stochastically bounded. Furthermore, a larger AR(1) coefficient $\rho$ tends to worsen the performance of $\hat{k}_{BKW}$, while it does not have effect on our QML estimator. As the sample size increases, the performance of $\hat{k}_{MS}$ improves, but it consistently underperforms QML in terms of RMSEs and MAEs.
Table 2: Simulated root mean squared errors (RMSEs) and mean absolute errors (MAEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, under DGP 1.B.
$N,T$ $(\hat{k}_{1},\hat{k}_{2})_{QML}$ $(\hat{k}_{1},\hat{k}_{2})_{BKW}$ $(\hat{k}_{1},\hat{k}_{2})_{MS}$
RMSE MAE RMSE MAE RMSE MAE
$\rho=0$ $\alpha=0$ $\beta=0$
100,100 (0.585,0.587) (0.238,0.220) (10.278,10.950) (4.934,5.411) (13.527,16.188) (9.112,11.066)
100,300 (0.492,0.542) (0.194,0.216) (3.754,2.906) (1.323,1.283) (38.707,40.302) (23.927,27.641)
300,300 (0.366,0.355) (0.114,0.108) (3.459,3.672) (1.275,1.251) (32.591,32.323) (19.131,18.664)
300,600 (0.354,0.372) (0.107,0.118) (1.899,1.991) (0.886,0.880) (30.847,33.765) (9.957,9.870)
600,600 (0.290,0.293) (0.080,0.076) (1.967,1.943) (0.864,0.864) (8.324,17.133) (1.518,2.590)
$\rho=0.7$ $\alpha=0$ $\beta=0$
100,100 (0.575,0.495) (0.199,0.173) (20.072,19.915) (14.724,14.257) (11.472,12.697) (7.487,7.809)
100,300 (0.504,0.427) (0.180,0.136) (30.374,30.487) (13.404,12.818) (31.692,31.232) (17.584,17.151)
300,300 (0.316,0.363) (0.088,0.106) (27.303,27.966) (11.891,11.699) (21.725,23.120) (9.920,8.905)
300,600 (0.315,0.305) (0.081,0.075) (15.284,15.701) (4.036,3.498) (9.122,14.798) (1.736,2.338)
600,600 (0.221,0.261) (0.041,0.060) (22.984,16.870) (5.098,3.842) (1.162,2.937) (0.662,0.671)
$\rho=0$ $\alpha=0.3$ $\beta=0$
100,100 (0.636,0.578) (0.268,0.232) (11.187,10.698) (5.679,5.243) (14.503,16.041) (10.362,10.710)
100,300 (0.551,0.611) (0.210,0.247) (4.328,4.518) (1.482,1.537) (34.096,33.528) (22.185,21.263)
300,300 (0.417,0.356) (0.130,0.111) (2.625,2.818) (1.122,1.237) (33.489,27.816) (19.837,15.085)
300,600 (0.370,0.369) (0.117,0.110) (2.049,1.685) (0.922,0.757) (31.595,28.276) (9.600,8.324)
600,600 (0.329,0.313) (0.094,0.088) (2.127,1.826) (0.937,0.825) (11.666,6.868) (2.513,1.230)
$\rho=0$ $\alpha=0$ $\beta=0.3$
100,100 (0.576,0.680) (0.240,0.282) (10.678,11.051) (5.083,5.397) (14.617,16.179) (10.349,10.967)
100,300 (0.624,0.551) (0.249,0.225) (3.181,2.893) (1.410,1.323) (36.529,31.525) (23.806,19.872)
300,300 (0.435,0.329) (0.143,0.096) (2.696,2.752) (1.227,1.146) (28.129,31.622) (16.158,16.622)
300,600 (0.362,0.363) (0.113,0.116) (2.100,1.877) (0.939,0.780) (31.471,37.708) (10.541,10.886)
600,600 (0.311,0.300) (0.087,0.082) (2.034,1.865) (0.898,0.833) (13.322,6.358) (2.953,1.107)
$\rho=0.7$ $\alpha=0.3$ $\beta=0.3$
100,100 (0.531,0.577) (0.192,0.179) (21.127,19.862) (15.306,14.535) (12.497,13.936) (8.482,8.872)
100,300 (0.504,0.427) (0.180,0.136) (30.374,30.487) (13.404,12.818) (31.692,31.232) (17.584,17.151)
300,300 (0.344,0.336) (0.082,0.089) (30.739,26.776) (12.881,11.394) (23.455,21.124) (10.231,7.694)
300,600 (0.295,0.321) (0.077,0.081) (16.831,17.168) (4.216,3.952) (5.083,6.411) (1.088,1.197)
600,600 (0.221,0.307) (0.049,0.074) (15.492,18.497) (3.935,4.216) (3.111,3.017) (0.836,0.587)
@@
DGP 1.C We set $\lambda_{i}$ to be i.i.d. $N(0,\frac{1}{r}I_{r})$ across $i$ and $\Lambda=(\lambda_{1},\cdots,\lambda_{N})^{{}^{\prime}}$. In the first regime, $\Lambda_{1}=(\lambda_{1,1},\cdots,\lambda_{N,1})^{{}^{\prime}}=\Lambda B_{1}$ with $B_{1}=[1,0,0;0,1,0;0,0,1]$. In the second regime, $\Lambda_{2}=(\lambda_{1,2},\cdots,\lambda_{N,2})^{{}^{\prime}}=\Lambda B_{2}$ with $B_{2}=[1,0,0;0,1;0,0,0]$. In the third regime, $\Lambda_{3}=(\lambda_{1,3},\cdots,\lambda_{N,3})^{{}^{\prime}}=\Lambda B_{3}$ with $B_{3}=[0,0,0;0,0,0;0,0,1]$. The number of pseudo-factors for the full sample is $r=3$, with the numbers of effective factors in the three regimes being 3, 2, and 1, respectively. Table 3 lists the RMSEs and MAEs of three estimators. In the first regime, all three factors are present. In the second regime, only the first two factors are active; and in the third regime, only the third factor is active.
In this DGP, both breaks are of Type 1. Specifically, the first breakpoint represents a scenario where some factors disappear after the break, satisfying the condition $r_{j,j+1}=\max\{r_{j},r_{j+1}\}$. From Table 3, we observe that when factors are correlated ($\rho=0.7$), the performance of the BKW method significantly deteriorates. The MS method also appears to struggle in this case, as both the RMSEs and MAEs of its first estimated breakpoint increase with sample size. This is attributed to the MS method’s assumption of a constant number of factors throughout the sample. Nevertheless, $\hat{k}_{QML}$ consistently demonstrates superior performance, with its RMSEs and MAEs tending towards zero as the sample size increases.

### v3

@@ 5 Simulation
where $u_{t}=(u_{t,1},\cdots,u_{t,r_{0}})^{{}^{\prime}}$ is i.i.d. $N(0,I_{r_{0}})$ for $t=2,\cdots,T$ and $f_{1}=(f_{1,1},\cdots,f_{1,r_{0}})^{{}^{\prime}}$ is drawn from $N(0,\frac{1}{1-\rho^{2}}I_{r_{0}})$. The scalar $\rho$ captures the serial correlation of factors, and the idiosyncratic errors are generated by
@@
where $v_{t}=(v_{1,t},\cdots,v_{N,t})^{{}^{\prime}}$ is i.i.d. $N(0,\Omega)$ for $t=2,\cdots,T$ and $e_{1}=(e_{1,1},\cdots,e_{N,1})^{{}^{\prime}}$ is $N(0,\frac{1}{(1-\alpha^{2})}\Omega)$. The scalar $\alpha$ captures the serial correlation of the idiosyncratic errors, and $\Omega$ is generated as $\Omega_{ij}=\beta^{|i-j|}$ so that $\beta$ captures the degree of cross-sectional dependence of the idiosyncratic errors. In addition, $u_{t}$ and $v_{t}$ are mutually independent for all values of $t$. We consider the following DGPs for factor loadings and investigate the performance of the QML estimator for the different types of breaks discussed in Section 3.
@@
DGP 1.A In this setup, we adopt the DGP from MS. Each of the three regimes has $r_{0}$ effective factors, i.e., $r_{1}=r_{2}=r_{3}=3$. The factor loading matrices on the effective factors are generated as follows:
@@
The $\Lambda$ matrix in (17) can be written as $\Lambda\;=\;\bigl[\breve{\Lambda}_{1},\breve{\Lambda}_{2},\breve{\Lambda}_{3}\bigr]_{N\times 3r_{0}}$. We define the three $3r_{0}\times 3r_{0}$ block-selector matrices
$B_{1}=\operatorname{diag}\!\bigl(I_{r_{0}},\,0_{r_{0}\times r_{0}},\,0_{r_{0}\times r_{0}}\bigr),\quad B_{2}=\operatorname{diag}\!\bigl(0_{r_{0}\times r_{0}},\,I_{r_{0}},\,0_{r_{0}\times r_{0}}\bigr),\quad B_{3}=\operatorname{diag}\!\bigl(0_{r_{0}\times r_{0}},\,0_{r_{0}\times r_{0}},\,I_{r_{0}}\bigr),$
@@
so that in regime $j$ we set $\Lambda_{j}=\Lambda\,B_{j}$, which yields three effective factors per regime and a total of $r=3r_{0}=9$ pseudo-factors over the full sample. This corresponds to a Type 1 breakpoint scenario with $r_{j,j+1}=r_{j}+r_{j+1}$, where loadings in adjacent regimes are independent. By varying $b$ we can control the magnitude of the break.
@@
DGP 1.B We set $\lambda_{i}$ to be i.i.d. $N(0,\frac{1}{r}I_{r})$ across $i$, and define $\Lambda=(\lambda_{1},\ldots,\lambda_{N})^{\prime}$. In the first regime, $\Lambda_{1}=(\lambda_{1,1},\cdots,\lambda_{N,1})^{{}^{\prime}}=\Lambda B_{1}$ with $B_{1}=[1,0,0;0,1,0;0,0,0]$. In the second regime, $\Lambda_{2}=(\lambda_{1,2},\cdots,\lambda_{N,2})^{{}^{\prime}}=\Lambda B_{2}$ with $B_{2}=[1,0,0;0,0,0;0,0,1]$. In the third regime, $\Lambda_{3}=(\lambda_{1,3},\cdots,\lambda_{N,3})^{{}^{\prime}}=\Lambda B_{3}$ with $B_{3}=[0,0,0;0,1,0;0,0,1]$. For the full sample, the number of pseudo-factors is $r=3$. In each of the three regimes, the number of effective factors is 2, as $\text{rank}(B_{1})=\text{rank}(B_{2})=\text{rank}(B_{3})=2$. This setting corresponds to a Type 1 break in Section 3 with $r_{j,j+1}>\max\{r_{j},r_{j+1}\}$.
@@
DGP 1.C The generation of $\Lambda$ remains the same as in DGP 1.B. For $B_{j}$, we set $B_{1}=[1,0,0;0,1,0;0,0,1]$, $B_{2}=[1,0,0;0,1,0;0,0,0]$ and $B_{3}=[0,0,0;0,0,0;0,0,1]$. The number of pseudo-factors for the full sample is $r=3$, with the numbers of effective factors in the three regimes being 3, 2, and 1, respectively. In the first regime, all three factors are present. In the second regime, only the first two factors are present and in the third regime, only the third factor is present.
DGP 1.D The generation of $\Lambda$ remains the same as in DGP 1.B. Let $\Lambda_{1}=\Lambda$, $\Lambda_{2}=2\Lambda$ and $\Lambda_{3}=\Lambda$. In this case, the number of factors remains constant across regimes, and the changes in factor loading are purely proportional, corresponding to a setting with two Type 2 breaks.
DGP 1.E In this setup, we consider a case involving both types of changes. The generation of $\Lambda$ remains the same as in DGP 1.B. We set $B_{1}=[1,0,0;0,1,0;0,0,0]$, $B_{2}=[2,x,y;0,2,z;0,0,0]$ where $x$, $y$, $z$ are i.i.d. $N(0,1)$ and $B_{3}=[0,0,0;0,0,0;0,0,1]$. By construction, there exists a nonsingular matrix $\bf R$} such that $B_{1}{\bf R}=B_{2}$, so the first breakpoint is Type 2 with $r=3>r_{1,2}=r_{1}=r_{2}=2$. The second breakpoint, however, is a Type 1 change and is responsible for the rank deficiency of the pseudo-factors in the first two regimes.
Since the computation of $\hat{k}_{MS}$ requires the number of effective factors ($r_{0}$), we use $\hat{r}=r_{0}$ for it. Moreover, as $\hat{k}_{QML}$ and $\hat{k}_{BKW}$ require the number of pseudo-factors ($r$), we set $\hat{r}=r$ for them. This allows us to focus on the estimation errors of the breakpoint estimators themselves, rather than errors resulting from misspecifying the number of factors.
Table 1 reports the RMSEs of the estimated break dates under DGP 1.A–1.E for $(\rho,\alpha,\beta)=(0.7,0.3,0.3)$. For DGP 1.A (both $b=1$ and $b=0$), the QML estimator yields very small RMSE in small samples, and the RMSE quickly decreases as $(N,T)$ increase; BKW and MS also improve with sample size but are uniformly less accurate than QML. In DGP 1.B, QML again achieves the smallest RMSE and continues to improve with $(N,T)$, whereas BKW exhibits non-vanishing (stochastically bounded) errors as the sample size grows; MS improves but remains less accurate than QML. In DGP 1.C, where factors disappear and the factor dimension changes across regimes, both competing methods deteriorate—especially MS, which relies on a constant factor dimension—while QML remains stable and its RMSE decreases with $(N,T)$. For DGP 1.D (rotational change), QML delivers small and bounded estimation errors, consistent with Theorem 2(ii), and it clearly outperforms BKW; MS again performs poorly and does not improve with sample size. Finally, for the mixed design in DGP 1.E (Type 2 followed by Type 1), QML yields bounded RMSE for the Type 2 breakpoint and near-zero RMSE for the Type 1 breakpoint, whereas both BKW and MS have substantially larger errors. Overall, the simulation evidence is consistent with Theorem 2: QML is consistent under Type 1 changes and attains $O_{p}(1)$ estimation errors under Type 2 changes. Additional results for other values of $(\rho,\alpha,\beta)$ are reported in Appendix D of the supplementary material.